Exercise 14.40 - Enhanced - with Feedback 10 of 32 > A Review | Constants Periodic Table The following reaction is first ordeExercise 14.40 - Enhanced - with Feedback 10 of 32 > Review Constants Periodic Table The following reaction is first order inWrite a rate law for the reaction. Write a rate law for the reaction. Rate=k Rate=k[A] Rate=k[A]2 Rate=k[A]3
Answer #1

it is given to assume rate constant k to be same in all case.

numerical value is same but it unit is different for different order.

k= rate/[A]^n

=(Ms-1)/(M^n)

you will get the unit for different order.

n represent order.

M represent Molarity.

s represent time.

Soln k = 0.033s- knit of k is set thing it is 1st order. Part A Rate = K[AL Rate = 0.053 (5.7x10-2) ---300 81031 Given K = 0.